OpenStudy (anonymous):

find 60% of 35. write 46% as a decimal. write 2/5 as a percent. solve eac proportion 7/11=x/33 12/20=15/x please show work

OpenStudy (anonymous):

I'll do the first three 60% of 35 = 21 46% = .46 2/5 = 40% Done on a calculator

OpenStudy (anonymous):

35 times 0.6 = 21

OpenStudy (anonymous):

46 = 46/100 = 0.46

OpenStudy (anonymous):

how do u figue it

OpenStudy (anonymous):

2/5 = 40/100 = 40%

OpenStudy (anonymous):

did that help a bit? :)

OpenStudy (anonymous):

7/11 = x/33 Cross multiply 11x = 231 divide by 11 x = 21

OpenStudy (anonymous):

7/11=x/33 multiply by 3 21/33=x/33 x=21

OpenStudy (anonymous):

how did u do the first one do u just multiply the 2

OpenStudy (anonymous):

Which one are you talking about?

OpenStudy (anonymous):

for the 2/5 part what do u do

OpenStudy (anonymous):

percentages can be written as fractions over 100, so make 2/5 a fraction over 100: 2/5=40/100=40%

OpenStudy (anonymous):

do you get it? :)

OpenStudy (anonymous):

kinda trying to get it to work on the caculator

OpenStudy (anonymous):

ty

OpenStudy (anonymous):

no problem :)